Calculate the standard enthalpy change for the reaction
2Al(s)+〖Fe〗_2 O_3 (s)→2Fe(s)+〖Al〗_2 O_3 (s)
given that
2Al(s)+3/2 O_2 (g)→〖Al〗_2 O_3 (s) ∆H_rxn^°=-1669.8 kJ/mol
2Fe(s)+3/2 O_2 (g)→〖Fe〗_2 O_3 (s) ∆H_rxn^°=-822.2 kJ/mol
1
Expert's answer
2020-09-21T06:30:17-0400

given that

2Al(s)+3/2 O2 (g)→Al2 O3 (s) ∆H_rxn°=-1669.8 kJ/mol................................1.


2Fe(s)+3/2 O2 (g)→Fe2O3(s) ∆H_rxn°=-822.2 kJ/mol ....................................2.


For Reaction '

2Al(s)+Fe2O3(s)→2Fe(s)+Al2O3 (s)


∆H_rxn° = eqn 1 - eqn 2 .


= -1669.8 - ( - 822.2 ) kj / mole . = -847.6 kj / mol .


 standard enthalpy change for the reaction = - 847.6 kj / mole .


2Al(s)+ Fe2 O3 (s)→2Fe(s) + Al2 O3 (s)